OpenStudy (anonymous):

true or false : (b^2)^2.4=b^4.4 ???

Directrix (directrix):

What do you think?

OpenStudy (anonymous):

false

Directrix (directrix):

(b^2)^2.4 = b ^ 2*2.4 = b ^ 4.8

OpenStudy (anonymous):

true ???

Directrix (directrix):

I agree with blanca --> False. Here's the original question: (b^2)^2.4=b^4.4 And, cranking out the problem gave (b^2)^2.4 = b ^ 2*2.4 = b ^ 4.8 b^4.4 is not equal to b ^ 4.8 so I say "False" as did you.

OpenStudy (anonymous):

false (b^2)^2.4 = b^4.4 =b^4+b^4